# Increasing the Utility of Synthetic Images through Chamfer Guidance

### v1

@@ 2 Related work
While most modern text-to-image diffusion models utilize classifier-free guidance [29], there has been increasing exploration of alternative guidance methods to further improve the sampling process. A subset of works focus on improving image quality with methods including modifications to classifier-free guidance to reduce over-saturation, such as Adaptive Projected Guidance (APG) [48], utilization of blurring [32, 31] or perturbations [2], and employment of a smaller, less-trained version of the model for guidance [34]. Other works explore guidance methods to improve the diversity of generated images, e.g.,  Condition-Annealed Diffusion Sampler (CADS) [47] anneals the conditioning signal by incorporating scheduled, monotonically decreasing Gaussian noise into the conditioning vector during the diffusion chain to increase diversity of the generated samples, or c-VSG, which guides the denoising process to increase the diversity of a sample compared to images previously generated with the same prompt [24]. Still other works focus on safety-related guidance by reducing harmful representations, such as using classifier guidance to guide generations away from inappropriate content [9, 49] or “forgetting” concepts by zeroing out their cross-attention scores [63]. However, these methods tend to focus on generation quality/diversity/safety as ends in themselves, rather than as means to increasing a synthetic image corpus’ utility in training downstream models.
@@ 4.3.1 Quantitative Results
Table 3: Validation accuracy on real images of classifiers trained with real data only, synthetic data only, and a mix of synthetic data and ImageNet training data for ImagetNet-1k classes. All models are trained for 200k iterations.
@@ 4.4 Training downstream image classifiers on synthetic data
We also test on ImageNet-V2 [44], ImageNet-Sketch [60], ImageNet-R [26], and ImageNet-A [27] to measure out-of-distribution (OOD) generalization. Models trained with data generated by our Chamfer guidance exhibit strong performance on these out-of-distribution datasets, surpassing models trained on only real data on all variants, given the same number of real images. For ImageNet-Sketch, LDM3.5M with our Chamfer guidance surpasses the performance of a classifier trained on the full (1.3M samples) ImageNet-1k. Our Chamfer guidance always obtains substantial gains over the default sampling of LDM1.5 and LDM3.5M across all ImageNet variants, up to $16\%$ when using synthetic data only for ImageNet-Sketch on LDM3.5M. Interestingly, on ImageNet-R and ImageNet-A, adding real data to our synthetic ones harms the model’s performance. In these domains, which differ significantly from ImageNet-1k, the gain resulting from synthetic data is less pronounced w.r.t. ImageNet-V2 and ImageNet-Sketch. R(endition) and A(dversarial) represent more challenging domains where more natural-looking images bring reduced benefit. Nevertheless, our Chamfer guidance still surpasses their real-only counterparts.
These results confirms the downstream utility of our generated samples for OOD tasks as well. Similarly to the in-domain results, we observe how our Chamfer guidance can effectively close the performance gap between LDM1.5 and LDM3.5M, and for ImageNet-Sketch, the performance of LDM3.5M is superior to LDM1.5.
@@ 4.5 Ablations
We conduct our ablation studies on LDM1.5, using the ImageNet-1k dataset. The goal is to understand the robustness of our Chamfer Guidance to the most relevant hyperparameters, i.e., the CFG scale $\omega$ from Equation˜2 (shown in Table˜4) and the strength of the Chamfer Guidance $\gamma$ from Equation˜6 (shown in the Appendix).
@@ 5 Conclusion
We introduced Chamfer Guidance, a novel training-free approach for improving the utility of synthetic data generated by conditional image models. By leveraging a small set of exemplar real images to guide the generation process, our method effectively balances quality and diversity while addressing distribution shifts between synthetic and real data distributions. Experimental results on ImageNet-1k and geo-diversity benchmarks demonstrate that Chamfer Guidance achieves state-of-the-art performance, and scales its effectiveness as the number of exemplars increases. Additionally, our synthetic data can be used in downstram applications as a synthetic data source. Furthermore, our approach eliminates the computational overhead associated with CFG, reducing computational requirements while maintaining superior performance in terms of quality and diversity. These contributions advance the field of synthetic data generation and demonstrate that thoughtfully designed guidance mechanisms can unlock the full potential of conditional image generative models for downstream machine learning applications.
Limitations. We build our evaluation of reference-based diversity and quality on automated metrics, which inherit intrinsic biases, e.g., the distribution of the dataset we compare against, or the reliance on pre-existing feature extractors, which might or might not capture subtle differences. We rely on previous works to provide the most accurate metrics as possible, although these metrics are statistical aggregators and do not account for individual preferences.
@@ Appendix B On the evaluation metrics
Appendix C Geographic diversity baselines
@@
Appendix D Additional geographic diversity results
@@
Appendix E Additional ablation
@@
Appendix F Additional object-centric results
In Table˜10 we show more results of base LDM1.5 and LDM3.5M sampling, and additional set of parameters for CADS [47] and APG [48]. These results extend those presented in Table˜1, and show how these reference-free approaches cannot significantly increase the fidelity and diversity of the samples, and that applying APG on a small $\omega$ brings moderate improvement over the best base sampling, for both LDM1.5 and LDM3.5M.
@@ Appendix F Additional object-centric results
Appendix G Additional classification results
We complement our downstream utility analysis by evaluating the downstream utility of images generated using our Chamfer Guidance. We follow the “static” ImageNet-100 setup introduced in [4], and generate a dataset of $130,000$ synthetic images using our Chamfer Guidance, using $k=32$ real exemplar images from the training set. Each synthetic image is generated with a simple prompt (class name). We report the accuracy of a ViT-B [15] classifier trained on this synthetic data and tested on real validation data. In Table˜11 we compare to the base performance of several LDMs, and observe substantial gains when using our Chamfer Guidance. Interestingly, the initial performance gap of 6.5 points between LDM1.5 and LDM3.5M (presumably because reduced sample diversity in LDM3.5M) is reduced to 2.4 points when using Chamfer Guidance. These results confirm the ability of Chamfer Guidance to improve downstream utility of synthetic data.
@@ Appendix G Additional classification results
Appendix H Licenses
We report the licenses for datasets and models used in Table˜12.
@@ Appendix H Licenses
Appendix I Additional qualitative results
Figure˜3 shows LDM1.5 with $\omega=1.0$ without and with our Chamfer guidance, and $\omega=2.0$ with our Chamfer guidance. LDM1.5 does not have a “strong” conditional only model ($\omega=1.0$), as quantitatively highlighted in Table˜10, which exhibits poor coverage and density, and a high FID. Qualitatively, we observe bad-looking images that often lack the correct subject, or present a distorted one. Our Chamfer guidance, even when applied only the conditional model, can effectively recover the correct texture, shape, and proportions of the subject. For LDM1.5, the best results are obtained by applying our Chamfer guidance on top of a low $\omega$ value, e.g.2.02.0.
@@ Appendix I Additional qualitative results
Figure˜5 qualitatively shows the effect of increasing the number of exemplar images $k$ when using our Chamfer guidance. We can see a greater diversity in both subjects and background when using a higher number for $k$, in agreement with the quantitative results presented in Figure˜2.
In Figure˜6 we show the comparison of the geographic diversity benchmark, in particular the GeoDE dataset. We show how base generations using LDM1.5 introduce stereotypical elements when using the geographical indicator, e.g., dilapidated cars, or rusty pans. c-VSG [24] partially mitigates the issue, but it comes at the expense of quality, generating oversaturated images that do not resemble the dataset. Our Chamfer guidance effectively mitigates both issues, by generating natural-looking images without stereotypical elements, due to the guidance of real exemplars.

### v2

@@ 2 Related work
While most modern text-to-image diffusion models utilize classifier-free guidance [30], there has been increasing exploration of alternative guidance methods to further improve the sampling process. A subset of works focus on improving image quality with methods including modifications to classifier-free guidance to reduce over-saturation, such as Adaptive Projected Guidance (APG) [50], utilization of blurring [33, 32] or perturbations [2], and employment of a smaller, less-trained version of the model for guidance [35]. Other works explore guidance methods to improve the diversity of generated images, e.g.,  Condition-Annealed Diffusion Sampler (CADS) [49] anneals the conditioning signal by incorporating scheduled, monotonically decreasing Gaussian noise into the conditioning vector during the diffusion chain to increase diversity of the generated samples, while Limited Interval guidance [37] removes the conditioning signal in a certain interval to improve ungrounded diversity. Particle Guidance [11] overcomes the common assumption of independent samples in a batched generation, and proposes an extension of sampling where a joint-particle time-evolving potential enforces diversity. c-VSG, which guides the denoising process to increase the diversity of a sample compared to images previously generated with the same prompt [25]. Still other works focus on safety-related guidance by reducing harmful representations, such as using classifier guidance to guide generations away from inappropriate content [9, 51] or “forgetting” concepts by zeroing out their cross-attention scores [65]. However, these methods tend to focus on generation quality/diversity/safety as ends in themselves, rather than as means to increasing a synthetic image corpus’ utility in training downstream models.
@@ 4.3.1 Quantitative Results
Table 3: Validation accuracy on real images of classifiers trained with real data, synthetic data, and a mix of synthetic and training data for ImagetNet-1k. All models are trained for 200k iterations.
@@ 4.4 Training downstream image classifiers on synthetic data
We also test on ImageNet-V2 [46], ImageNet-Sketch [62], ImageNet-R [27], and ImageNet-A [28] to measure out-of-distribution (OOD) generalization. Models trained with data generated by our Chamfer guidance exhibit strong performance on these out-of-distribution datasets, surpassing models trained on only real data on all variants, given the same number of real images. For ImageNet-Sketch, LDM3.5M with our Chamfer guidance surpasses the performance of a classifier trained on the full (1.3M samples) ImageNet-1k. Our Chamfer guidance always obtains substantial gains over the default sampling of LDM1.5 and LDM3.5M across all ImageNet variants, up to $16\%$ when using synthetic data only for ImageNet-Sketch on LDM3.5M. Interestingly, on ImageNet-R and ImageNet-A, adding real data to our synthetic ones harms the model’s performance. In these domains, which differ significantly from ImageNet-1k, the gain resulting from synthetic data is less pronounced w.r.t. ImageNet-V2 and ImageNet-Sketch. R(endition) and A(dversarial) represent more challenging domains where more natural-looking images bring reduced benefit. Nevertheless, our Chamfer guidance still surpasses their real-only counterparts. These results confirms the downstream utility of our generated samples for OOD tasks as well. Similarly to the in-domain results, we observe how our Chamfer guidance can effectively close the performance gap between LDM1.5 and LDM3.5M, and for ImageNet-Sketch, the performance of LDM3.5M is superior to LDM1.5.
@@ 4.5 Ablations
[…]
@@ 5 Conclusion
We introduced Chamfer Guidance, a novel training-free approach to improve the utility of synthetic data from conditional image models. By leveraging a small set of exemplar real images to guide the generation process, our method balances quality and diversity while addressing distribution shifts between synthetic and real data. Experimental results on ImageNet-1k and geo-diversity benchmarks show Chamfer Guidance achieves state-of-the-art performance and scales its effectiveness as exemplars increase. Additionally, our synthetic data can be used in downstream applications. Furthermore, our approach eliminates the computational overhead of CFG, reducing computational requirements while maintaining superior performance in quality and diversity. These contributions advance synthetic data generation and demonstrate that thoughtfully designed guidance can unlock the full potential of conditional image generative models for downstream applications.
Limitations. Our evaluation of reference-based diversity and quality is built on automated metrics, which inherit intrinsic biases, e.g., , the distribution of the dataset we compare against, or the reliance on pre-existing feature extractors that might not capture subtle differences. While we rely on previous works to provide the most accurate metrics possible, these are statistical aggregators and do not account for individual preferences. Currently, our approach is designed for class-conditional models and does not directly support text-to-image generation. This limitation opens up several exciting avenues for future work. To extend our method to text-to-image models, we envision a retrieval-based pipeline. First, a large text-image dataset would be embedded into semantic vectors to create an offline retrieval database. Then, at inference time, for a user’s text query, the top-k relevant images would be retrieved to serve as exemplars for our guidance. Another interesting extension would be a zero-shot, data-free pipeline. We could envision a self-bootstrapping technique that first generates initial candidate images for a class, then automatically selects a diverse subset maximizing coverage (or “diameter”) in a robust feature space, and finally uses this synthetically-generated set as the guidance exemplars.
@@
Acknowledgments. This work was sponsored by the project FAIR Future AI Research (PE00000013), funded by NextGeneration EU.
@@ Appendix B On the evaluation metrics
Appendix C User Study
While the primary focus of our study is the utility of the generated data for representation and as a training source for downstream tasks, which is confirmed by our extensive experimental evaluation with downstream classifier training, we also evaluate human perception of our generated images.
To this end, we conduct a small-scale user study to complement our quantitative findings. We collected 965 data points from more than twenty anonymous annotators. In this study, users were presented with samples generated from prompts based on the ImageNet dataset. Their task was to choose their preferred generation in a side-by-side comparison between images from the base LDM3.5M model and images generated with our Chamfer Guidance applied to the same model. Users were also presented with real images from the dataset to ground their evaluation in real-world quality and coherence, as we show in Figure˜3.
The results showed a strong preference for our method: images generated with our Chamfer Guidance were preferred in $\mathbf{92\%\pm 2\%}$ of the cases. This suggests that the automatic evaluation of quantitative improvements in downstream utility also correlates with enhanced human-perceived quality and fidelity to the target concept distribution.
Figure 3: Example of user study question.
Appendix D Geographic diversity baselines
@@
Appendix E Additional geographic diversity results
@@
Appendix F Additional ablation
@@
Feature extractor ablation.
To further validate the choice of DINOv2, we conducted a preliminary empirical study comparing its performance as a feature extractor for performing Chamfer guidance against CLIP on the GeoDE dataset in an “object-centric” setting, e.g., , with prompts like "a photo of a car.". These experiments were run using LDM1.5.
Our findings were as follows:
  • DINOv2 consistently yielded the best performance, showing substantial gains in both diversity and fidelity. Similarity in DINOv2 space also correlates to human-perceived similarity slightly more than CLIP, as reported in [24].
  • CLIP improved upon the baselines, and in particular, coverage scales with the number of guiding samples. With higher $k$ we observe reduced marginal improvements, and we deem this to the fact that the model tended to converge towards generating an average representation of the object. We hypothesize this is because CLIP’s pre-training is “concept-centric” (aligning images to general text concepts), whereas DINOv2’s is “instance-centric” due to its self-supervised training, making it better at preserving the unique features of a specific reference image.
We report these results in Table˜10. This analysis confirms that DINOv2 is the most effective choice for our method, with CLIP being an alternative.
Table 10: Comparison of feature extractors across different $k$ values. DINOv2 shows strong improvements in $F_{1}$, precision, and coverage compared to LDM1.5 and CLIP, while also achieving lower FID.
Feature Extractor $k$ $\mathbf{F_{1}(P,C)}$ $\uparrow$ Precision $\uparrow$ Coverage $\uparrow$ Density $\uparrow$ Recall $\uparrow$ FDD $\downarrow$ FID $\downarrow$
LDM1.5 $\omega=1.0$ $0.2334$ $0.4363$ $0.1593$ $0.1731$ 0.6135 $684.81$ $35.27$
LDM1.5 $\omega=2.0$ $0.3277$ $0.5433$ $0.2346$ $0.2647$ $0.4859$ $524.59$ $24.04$
LDM1.5 $\omega=7.5$ $0.2960$ $0.6222$ $0.1942$ $0.3629$ $0.2025$ $693.31$ $42.60$
DINOv2 2 $0.4242$ $0.6354$ $0.3184$ $0.4027$ $0.4614$ $410.84$ $19.73$
DINOv2 4 $0.5313$ $0.8296$ $0.3907$ $0.8933$ $0.1501$ $368.27$ $19.08$
DINOv2 8 $0.6251$ $0.9010$ $0.4785$ $1.3575$ $0.0708$ $353.80$ $18.81$
DINOv2 16 0.7527 0.9461 0.6250 2.0309 $0.0547$ 323.11 18.44
CLIP 2 $0.3888$ $0.6367$ $0.2798$ $0.3987$ $0.3604$ $446.77$ $19.66$
CLIP 4 $0.4063$ $0.5951$ $0.3084$ $0.3501$ $0.4827$ $421.20$ $19.01$
CLIP 8 $0.4163$ $0.5974$ $0.3195$ $0.3567$ $0.5030$ $416.75$ $18.25$
CLIP 16 $0.4088$ $0.5757$ $0.3169$ $0.3180$ $0.5735$ $415.02$ $18.22$
Appendix G Additional object-centric results
In Table˜11 we show more results of base LDM1.5 and LDM3.5M sampling, and additional set of parameters for CADS [49] and APG [50]. These results extend those presented in Table˜1, and show how these reference-free approaches cannot significantly increase the fidelity and diversity of the samples, and that applying APG on a small $\omega$ brings moderate improvement over the best base sampling, for both LDM1.5 and LDM3.5M.
@@ Appendix G Additional object-centric results
Appendix H Additional classification results
We complement our downstream utility analysis by evaluating the downstream utility of images generated using our Chamfer Guidance. We follow the “static” ImageNet-100 setup introduced in [4], and generate a dataset of $130,000$ synthetic images using our Chamfer Guidance, using $k=32$ real exemplar images from the training set. Each synthetic image is generated with a simple prompt (class name). We report the accuracy of a ViT-B [16] classifier trained on this synthetic data and tested on real validation data. In Table˜12 we compare to the base performance of several LDMs, and observe substantial gains when using our Chamfer Guidance. Interestingly, the initial performance gap of 6.5 points between LDM1.5 and LDM3.5M (presumably because reduced sample diversity in LDM3.5M) is reduced to 2.4 points when using Chamfer Guidance. These results confirm the ability of Chamfer Guidance to improve downstream utility of synthetic data.
@@ Appendix H Additional classification results
Appendix I Licenses
We report the licenses for datasets and models used in Table˜13.
@@ Appendix I Licenses
Appendix J Additional qualitative results
Figure˜4 shows LDM1.5 with $\omega=1.0$ without and with our Chamfer guidance, and $\omega=2.0$ with our Chamfer guidance. LDM1.5 does not have a “strong” conditional only model ($\omega=1.0$), as quantitatively highlighted in Table˜11, which exhibits poor coverage and density, and a high FID. Qualitatively, we observe bad-looking images that often lack the correct subject, or present a distorted one. Our Chamfer guidance, even when applied only the conditional model, can effectively recover the correct texture, shape, and proportions of the subject. For LDM1.5, the best results are obtained by applying our Chamfer guidance on top of a low $\omega$ value, e.g.2.02.0.
@@
Figure˜6 qualitatively shows the effect of increasing the number of exemplar images $k$ when using our Chamfer guidance. We can see a greater diversity in both subjects and background when using a higher number for $k$, in agreement with the quantitative results presented in Figure˜2.
In Figure˜7 we show the comparison of the geographic diversity benchmark, in particular the GeoDE dataset. We show how base generations using LDM1.5 introduce stereotypical elements when using the geographical indicator, e.g., dilapidated cars, or rusty pans. c-VSG [25] partially mitigates the issue, but it comes at the expense of quality, generating oversaturated images that do not resemble the dataset. Our Chamfer guidance effectively mitigates both issues, by generating natural-looking images without stereotypical elements, due to the guidance of real exemplars.